OpenStudy (anonymous):

Evaluate v(w + 3) for v(w) = 2w - 1

OpenStudy (xapproachesinfinity):

How would you evaluate this f(2) given f(x)=2x-1? do you know how?

OpenStudy (xapproachesinfinity):

@battlefieldforlife

OpenStudy (anonymous):

v(w) = 2w - 1 Here see that v is given as a function of w. So, v(w+3) means that v is being expressed as a function of w+3. Thus we need to fill w+3 in place of w in the 1st equation. v(w+3) = 2(w+3) - 1 => v(w+3) = 2w + 6 - 1 = 2w + 5